### 19 Mar 2026 v1

# Comments on the Emergence of 4D Topological Amplitudes in M-Theory

Abstract
The M-theoretic Emergence Proposal claims that all of the terms in the low-energy action arise from quantum effects. After reviewing the current status of this proposal, we focus on four-dimensional compactifications with $N=2$ supersymmetry, where kinetic terms are encoded in topological string amplitudes, such as the prepotential $\mathcal{F}_{0}$. Evidence for the emergence of such terms was provided recently, where in particular it was shown that the classical cubic term in $\mathcal{F}_{0}$ can be obtained by integrating out the light towers of states in the M-theory limit, using a novel regularization of the infinite sum over Gopakumar/Vafa invariants. We address two issues that were left open. First, we show that the regularization can be equivalently performed in complex structure moduli space and in Kähler moduli space. Second, we extend the proposed regularization to the linear terms in the one-loop prepotential $\mathcal{F}_{1}$.
1 Introduction
One of the central aspects of the swampland program (see [1, 2, 3] for reviews) is the realization that string theory, or more generally Quantum Gravity (QG), admits infinite distance limits in its moduli space where infinite towers of states become exponentially light. This is captured by the swampland distance conjecture [4], which has been analyzed and refined during recent years. One essential refinement is the emergent string conjecture [5], which asserts that the lightest of such towers can only be of two types. These are either Kaluza-Klein (KK) modes of a decompactifying direction or the excitation modes of a weakly coupled, critical string, both meant in an appropriate duality frame.
In addition to the lightest towers appearing at infinite-distance limits, it was found that there exists a hierarchical pattern of all towers, namely that one can distinguish between light and heavy towers [6, 7], where the former ones are those with a typical mass scale not larger than the QG cut-off, which is taken to be the species scale [8, 9]. Similar to the structure of the weakly coupled fundamental string, the light towers can be considered the fundamental quantum degrees of freedom and the heavy ones as classical non-perturbative contributions. From this perspective, standard perturbative string theory is just the perturbative QG theory that arises in the small string coupling regime, i.e. an infinite distance limit for the dilaton. The perturbative states are the vibration modes of the string and upon compactification also its KK and winding modes. The non-perturbative states are the $NS5$-brane together with the various $D$-branes of the superstring theories, all with tensions larger than the species scale.
In an attempt to find a microscopic explanation to the swampland distance conjecture the so-called Emergence Proposal [10, 11, 12] was put forward (for subsequent work see e.g. [13, 14, 15]), which suggests that the kinetic terms in the low-energy effective action of (all) QG theories are emerging quantum mechanically from integrating out states below a certain UV scale. While initially less understood than some other swampland conjectures, it was subsequently realized that there exists a special limit in the QG moduli space where the conjecture seems to be realized in an even stronger sense [6, 7, 16] (for a complementary approach see [17, 18, 19, 20]). First, consider a compactification of type IIA string theory to $d$ dimensions on an internal space with radii $R_{i}$. The M-theory limit is obtained by scaling the type IIA string coupling and the radii to infinity in such a way that, in M-theory units, one has
[…]
where $R_{11}$ denotes the radius of the M-theory circle and $M_{*}$ the eleven-dimensional Planck scale. In this limit one takes the strong coupling regime of type IIA string theory, in which a new eleventh dimension opens up. At the same time, the size of the internal space in units of the eleven-dimensional Planck scale, as well as the $d$-dimensional Planck scale, are kept finite. This limit is called the M-theory limit [7] and is a decompactification to $d+1$ dimensions where the species scale is the higher dimensional Planck scale. Second, it was proposed that in this limit the full low-energy effective action arises from integrating out the light towers of states, namely those whose typical mass scales do not exceed the species scale. In other words, all interactions are generated through quantum effects, without any classical contribution (see [21] for a review).
The central challenge in providing evidence for this proposal is that we do not understand the full quantization of M-theory yet. Guided by the BFSS matrix model [22] (see [23] for a review), one realizes that this form of the Emergence Proposal is consistent with it. There, the fundamental degrees of freedom are matrices and a large scale commuting space-time can arise by solving the classical equations of motion, in particular by choosing diagonal matrices. However, interactions like gravity are known to be obtained from a quantum effect, namely from integrating out the fluctuations of matrices around the classical background.
However, for more involved compact backgrounds, a similar matrix model computation is not feasible with current techniques. Nonetheless, these difficulties can be bypassed by considering certain couplings in the effective action that are protected by supersymmetry and only receive contributions from $1/2$-BPS states. The latter are well understood and admit the usual geometric interpretation we are used to from string theory. Indeed, all current evidence for the M-theoretic Emergence Proposal are based on the evaluation of such $1/2$-BPS saturated couplings. Explicit checks were carried out by deriving these couplings from a one-loop Schwinger integral, where only the light, perturbative towers of particle-like states with masses below the species scale, namely the 11D Planck scale, are integrated out.
Let us briefly recall which consistency checks have been performed:
  • $\mathbf{R^{4}}$-term in supergravity theories with 32 supercharges: Arising from toroidal compactifications of type IIA/M-theory, the analysis builds on the pioneering work of Green-Gutperle-Vanhove [24], in which the couplings in ten and nine dimensions were obtained by integrating out the KK spectrum along the eleventh direction, i.e. bound states of $D0$-branes. In [25] tools were developed to regularize the appearing Schwinger integrals and shown to be consistent with string theory results. This was then applied to the evaluation of the $R^{4}$ Schwinger integral, where the rest of the light towers of states were integrated out. For dimensions $d\geq 7$ integrating out the full set of light towers leads to a perfect match with the exact amplitude. This analysis will be further refined in [26], where the various infinite distance limits as classified by the taxonomy rules [27] will be analyzed in a systematic manner. Importantly, only the M-theory limit from eq.(1) remains as the limit where the full $R^{4}$-term arises from integrating out the light modes.
  • $\mathbf{F^{4}}$-term in six-dimensional supergravity with 16 supercharges: This gauge coupling is obtained by compactifying type IIA on a $K3$ surface or the dual heterotic string on a $T^{4}$. Building on [28] and exploiting the duality at the special orbifold point $T^{4}/\mathbb{Z}_{2}$ in the $K3$ moduli space, [29] focuses on the $F^{4}$-coupling for a linear combination of 16 $U(1)$ gauge fields, each arising from the dimensional reduction of the Ramond - Ramond (RR) 3-form $C_{3}$ along a two-sphere. With a careful case-by-case study of the worldsheet instanton corrections on the heterotic side, it was possible to show that $D4$-brane contributions are indeed redundant, in the sense that they lead to mutually canceling terms, as their mass scale is above the species scale.
  • The topological string couplings ${\cal F}_{0}$ and ${\cal F}_{1}$ in four-dimensional $N=2$ supergravity (8 supercharges): These are related to F-terms in the effective action of type IIA compactified on a Calabi-Yau (CY) threefold. Note that ${\cal F}_{0}$ contains information about the second order supergravity action, namely about the gauge couplings and kinetic terms of the vector multiplet fields. Integrating out the light towers is precisely what has been done in the seminal work of Gopakumar/Vafa [30, 31], where the authors expressed the topological string amplitudes ${\cal F}_{g}$ for genus $g\geq 0$ in terms of a Schwinger integral, integrating out particle-like bound states of $D0$- and $D2$-branes. Although originally only obtaining the worldsheet instanton contributions, it was shown in [6] that for the non-compact resolved conifold also the classical polynomial contributions could be obtained by an appropriate regularization of the UV divergences. These results were the building blocks to generalize the computation to compact CY manifolds. In this case, the main challenge is a final regularization of the infinite sum over the so-called Gopakumar/Vafa (GV) invariants, which generically grow exponentially. This problem was approached in [32], where we provided a concrete recipe of how such a regularization procedure can be implemented. A key aspect of the regularization is that it involved taking a limit to certain singularities in the Kähler moduli space. However, for technical reasons, this limit was actually taken in the mirror dual complex structure moduli space.
In this article we focus the attention on this last setup of topological couplings ${\cal F}_{0}$ and ${\cal F}_{1}$ in four-dimensional $N=2$ supergravity. We first summarize the main idea of [32], which provided a regularization of the sum over all GV invariants for ${\cal F}_{0}$. Instead of revisiting in detail the computation presented there, this proceedings article focuses on two specific issues that emerged during discussions of this work. On the one hand we reconsider the regularization of ${\cal F}_{0}$, originally performed in the mirror complex structure moduli space, and confirm whether it yields the same result as taking the corresponding limit in the original Kähler moduli space. On the other hand, it is known that also ${\cal F}_{1}$ receives polynomial, in fact linear, contributions so a natural question is whether our regularization procedure can also be extended to this case. In both cases we approach these problems for the case of the Quintic.
2 Emergence of 4D $N=2$ topological couplings
It is known that the topological amplitudes ${\cal F}_{g}$ in type IIA compactifications on CY threefolds are $1/2$-BPS saturated. This yields $N=2$ supersymmetry in four dimensions, and the ${\cal F}_{g}$ only depend on the Kähler moduli. Since this is a well studied class of models, let us only very briefly recall some relevant facts.
2.1 Preliminaries
The corresponding vector-multiplet moduli space is spanned by real scalars $\tau_{i}$ which, together with the Kalb-Ramond axions $b_{i}$, define the complex Kähler moduli $t_{i}=b_{i}+i\tau_{i}$. The kinetic terms and gauge couplings of the vector-multiplets are determined by a holomorphic prepotential $\mathcal{F}_{0}(t)$. Due to supersymmetry, this is tree-level exact in type IIA and for a CY with a single Kähler modulus reads
$\mathcal{F}_{0}(t)=\frac{(2\pi i)^{3}}{g_{s}^{2}}\Big[\frac{1}{6}\kappa_{ttt}\,t^{3}-\frac{\zeta(3)}{2(2\pi i)^{3}}\ \chi+\frac{1}{(2\pi i)^{3}}\sum_{\beta\in H_{2}(X,\mathbb{Z})}\alpha_{0}^{\beta}\,{\rm Li}_{3}(e^{2\pi i\beta\cdot t})\Big]\,$, (2)
where $\kappa_{ttt}$ denotes the triple intersection number of the CY $X$, $\chi$ its Euler characteristic and $\alpha_{0}^{\beta}$ are genus zero GV invariants [30, 31] [^1] Note that mirror symmetry fixes the a priori ambiguous quadratic and linear terms in the prepotential, which we are ignoring here.. In addition there are higher genus contributions ${\cal F}_{g}$ to the prepotential, which are the coefficients of higher derivative corrections. Similarly to the triple intersection number in $\mathcal{F}_{0}$, the genus-one prepotential $\mathcal{F}_{1}$ contains a classical term at weak coupling, which for a one-parameter CY threefold reads
$\mathcal{F}_{1}(t)={2\pi i}\left(\frac{c_{2}}{24}t+\frac{1}{2\pi i}\sum_{\beta\in H_{2}(X,\mathbb{Z})}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right){\rm Li}_{1}(e^{2\pi i\beta\cdot t})\right)\,$, (3)
where $c_{2}$ is the second Chern class of the manifold and $\alpha_{1}^{\beta}$ are the genus one GV invariants [30, 31].
The M-theoretic Emergence Proposal asserts that the full expressions can be reproduced in the M-theory limit (1) from a single Schwinger integral. In this limit, the M-theory scaling corresponds to rescaling all type IIA Kähler moduli as $\tau_{i}\rightarrow\lambda\,\tau_{i}$ and the string coupling as $g_{s}\rightarrow\lambda^{3/2}g_{s}$, with $\lambda\rightarrow\infty$, while keeping the four-dimensional Planck scale fixed. In this regime, the $D0$-branes become the lightest states and determine the species scale $\Lambda_{\rm sp}\sim M_{\rm pl}/\lambda^{1/2}$, which lies at the threshold of the mass of $D2$-branes wrapping 2-cycles of the CY manifold. Expressed in M-theory units, the co-scaled limit keeps the radii of the CY manifold fixed when measured in units of the eleven-dimensional Planck scale $M_{*}$, while the M-theory circle radius and the Planck scale scale as $r_{11}\rightarrow\lambda\,r_{11}$ and $M_{*}\rightarrow M_{*}/\lambda^{1/2}$, respectively. In such a limit the species scale coincides with the five-dimensional Planck scale, which scales in the same way as $M_{*}$.
Integrating out the light towers of states with mass scale below or at the species scale we arrive at the prescription of Gopakumar/Vafa [30, 31], i.e. one has to integrate out wrapped $M2$-branes transverse with respect to the M-theory circle but with KK-momentum along it. The corresponding Schwinger integral for $\mathcal{F}_{0}$ reads
$\mathcal{F}_{0}(t)=\sum_{(\beta,n)\neq(0,0)}\alpha_{0}^{\beta}\int_{0}^{\infty}\frac{ds}{s^{3}}\;e^{sZ_{\beta,n}}\,$, (4)
where $Z_{\beta,n}=\frac{2\pi i}{g_{s}}(\beta\cdot t-n)$ is the central charge of the supersymmetry algebra. Similarly, for $\mathcal{F}_{1}$ Gopakumar/Vafa obtained
$\mathcal{F}_{1}(t)=\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\int_{0}^{\infty}\frac{ds}{s}e^{sZ_{\beta,n}}\,.$ (5)
We recall that, after employing a minimal subtraction of the UV divergence at $t\to 0$ and subsequently applying $\zeta$-function regularization, the complete $\mathcal{F}_{0}$ and $\mathcal{F}_{1}$ for the (non-compact) resolved conifold have been reproduced in [6]. This included also the cubic and linear term in $\mathcal{F}_{0}$ and $\mathcal{F}_{1}$, respectively. In that case, the sum over GV invariants in (4) and (5) was absent. However, for a general compact CY, the $D2$-$D0$ bound states wrap all 2-cycles in $H_{2}(X,\mathbb{Z})$ and therefore another infinite sum needs to be performed.
How such a regularization of the infinite sum over GV invariants can be performed was analyzed in [32], at least for the prepotential $\mathcal{F}_{0}$. It is easier to evaluate its third derivative with respect to $t$, giving the Yukawa coupling, for which the integral over $t$ can be performed explicitly
$Y_{ttt}=\frac{g_{s}^{2}}{(2\pi i)^{3}}\partial_{t}^{3}\mathcal{F}_{0}(t)=\sum_{\beta\neq 0}\alpha_{0}^{\beta}\,\beta^{3}\left(\frac{1}{2}+\frac{e^{2\pi i\beta t}}{1-e^{2\pi i\beta t}}\right)\,.$ (6)
Here the second term in the brackets is related to the non-perturbative terms in (2), i.e. the ${\rm Li}_{3}$-terms and the factor $1/2$ matches precisely the formal triple intersection number for the single two-cycle of the resolved conifold [33]. Note that for a compact CY, the first term still involves the infinite sum over all GV invariants, which was called the zero point Yukawa coupling in [32] and was denoted as $Y_{ttt}^{(0)}$. This object is the focus of our interest, as the Emergence Proposal implies that computing this quantity in the strongly coupled M-theory limit should give the triple intersection number of the CY. To appreciate the task, recall that the exact determination of $Y^{(0)}_{ttt}$ in a fully fledged microscopic description of M-theory would (likely) lead to a finite result right away. Since such a theory is lacking at present, the best one can try for is to find a properly working regularization scheme.
In the following we do not repeat the detailed derivation leading to this proposal and instead provide a heuristic motivation. For the complete computation we refer the reader to the original work [32]. Implementing the zeta function regularization $\sum_{k=1}^{\infty}1=\zeta(0)=-1/2$ one writes
$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{n}n^{3}=-\sum_{n,k=1}^{\infty}\alpha_{n}n^{3}=-\lim_{\tau\to 0}\sum_{n,k=1}^{\infty}\alpha_{n}n^{3}e^{-2\pi nk\tau}=-\lim_{\tau\to 0}\sum_{n=1}^{\infty}\alpha_{n}n^{3}\frac{q^{n}}{1-q^{n}}\end{split}$ (7)
with $q=\exp(-2\pi\tau)$. Here, one recognizes the last term as the instanton terms in (6). To evaluate that sum one needs to determine the GV invariants. Determining these invariants by direct counting is computationally out of reach in general; however, their values are known from mirror symmetry. Hence, setting for ease of notation $(2\pi i)^{3}/g_{s}^{2}=1$, we can express the last term in (7) as
$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{n}n^{3}=-\lim_{\tau\to 0}\sum_{n}^{\infty}\alpha_{n}n^{3}\frac{q^{n}}{1-q^{n}}=-\lim_{t\to 0}\left[(\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}-\kappa_{ttt})\right]\end{split}$ (8)
where we have identified $t=i\tau$ and $\mathcal{F}_{0}|_{\rm weak}$ denotes the prepotential that one determines by determining periods in the complex structure moduli and applying the mirror map. This formula now opens up the possibility to define the regularization of the zero-point Yukawa coupling by minimal subtraction of the divergent terms in the limit $t\to 0$ . Hence, we define the regularization
$Y_{ttt}^{(0)}=-\lim_{t\rightarrow 0}\left[(\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}-\kappa_{ttt})-{\rm Div}\right]\,$, (9)
where Div denotes all terms that diverge in the limit.
Nonetheless, there is still the issue that it involves taking the limit $t\rightarrow 0$. In fact, as already observed in [34], the exponential degeneracy of the GV invariants is correlated with the appearance of a divergence in the prepotential, not at zero but at a finite value $t=t_{c}$ determined by the location of the conifold singularity. Indeed, the point $t=0$ is not even part of the quantum moduli space. Moreover, $t_{c}$ is not the minimal value of $\mathrm{Im}(t)$, which is instead given at the Landau–Ginzburg point. The conifold point is nothing else than the quantum-corrected singular point in the middle of the classical phase diagram of a Gauged Linear Sigma Model (GLSM).
In this sense, the limit can be interpreted as approaching zero size, corresponding to the vanishing of the Fayet–Iliopoulos (FI) parameter, $\xi\to 0$. However, since it is known (see e.g. [35]) that due to quantum effects the location of the singularity changes, the most natural thing to do is to take the limit towards its quantum corrected value so that we define
$Y_{ttt}^{(0)}=-\lim_{t\rightarrow t_{c}}\left[(\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}-\kappa_{ttt})-{\rm Div}\right]\,.$ (10)
A few remarks about this prescription are in order:
  • Since the wanted Yukawa coupling $\kappa_{ttt}$ already explicitly appears on the right hand side, the procedure works if the limit of $\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}$ only involves divergent terms and terms that vanish in the $t\to t_{c}$ limit.
  • This formula is not trivial, as we are not taking the limit $t\to i\infty$ where of course $\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}$ reproduces $\kappa_{ttt}$. To remind the reader, here we propose a procedure that is valid in the M-theory limit, i.e. strong string coupling regime.
  • Clearly, to evaluate the limit one must know the form of the prepotential $\mathcal{F}_{0}$ close to the conifold point. In practice, this is most conveniently obtained on the mirror side by solving the Picard-Fuchs equations in a local coordinate of the conifold patch of the mirror CY and picking a symplectic basis of periods by analytic continuation from the large complex structure (LCS) patch. We quickly recall how this is done later. Finally, one can then rewrite it in the Kähler modulus by inverting the mirror map $t(u)$. How all these computations are carried out in detail is described in the original work [32].
Although in this proceeding article we focus on the Quintic, in [32] the regularization procedure was generalized to CY manifolds with $h^{1,1}=2$. In those cases, which codimension-two singular point the limit was taken to was crucial in determining whether the regularization procedure worked. Although this issue remains unresolved, we focus here on two simpler questions that were left open.
3 Regularization scheme for the prepotential ${\cal F}_{0}$
Since inverting the mirror map is technically challenging, in [32] the regularization was performed directly on the mirror side by subtracting the divergent terms in the complex structure modulus $u$. One may wonder whether, upon reexpressing the result in terms of the Kähler modulus $t$, some of the divergences subtracted in (10) could give rise to a finite constant contribution. Since we are interested in the original CY, if that was case the minimal subtraction performed in $u$ would not be sufficient and the regularization prescription would have to be adjusted.
3.1 Limit in complex structure moduli space
To show that this does not occur we revisit the simple example of the quintic and recall how the computation was carried out in the complex structure coordinate. Around the conifold point, $\psi=1$, it is convenient to introduce the local coordinate $u=1-\psi^{-5}$. Solving the Picard-Fuchs equations and analytically continuing the symplectic basis to the conifold patch gives the mirror map [32]
$t=\frac{X_{1}}{X_{0}}=t_{c}\sum_{m=0}^{\infty}(u\log u)^{m}c_{m}(u)\,$, (11)
where $t_{c}\approx i1.20812$ and $c_{m}(u)=\sum_{n=0}^{\infty}c_{(m,n)}u^{n}$ denote infinite series in $u$. Some low-order coefficients are listed in Table 1.
[…]
$m=0$ 11 $0.139626$ $0.074073$ $0.049758$
$m=1$ 0.0528989-0.0528989 0.045636-0.045636 0.0396285-0.0396285 0.0351159-0.0351159
Table 1: Coefficients $c_{(m,n)}$ of the mirror map.
Next, recall that the Yukawa coupling can be expressed as [34]
$\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}=\frac{1}{\omega_{0}^{2}}\kappa_{\psi\psi\psi}\frac{1}{(dt/d\psi)^{3}}$ (12)
with
$\kappa_{\psi\psi\psi}=\left(\frac{2\pi i}{5}\right)^{3}\frac{5\psi^{2}}{1-\psi^{5}}\,.$ (13)
Using $u=1-\psi^{-5}$ and (11), each factor admits an expansion in powers of $u$ and inverse powers of $\log u$. Specifically,
[…]
where again the $d_{n}(u)$ and $e_{n}(u)$ are infinite series in $u$. Importantly, we will later use the fact that
$d_{-1}(u)=\frac{d_{(-1,1)}}{u}+d_{(-1,0)}+\mathcal{O}(u)\,.$ (15)
Putting everything together and performing an expansion around $u\to 0$, we obtain
$\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}=\frac{1}{u\log^{3}u}\left(\sum_{n,k=0}^{\infty}\frac{(u\log u)^{n}}{\log^{k}u}a_{(n,k)}(u)\right)\ ,$ (16)
where $a_{(n,k)}(u)$ denotes an infinite series in $u$. The terms split into a divergent piece when $u\rightarrow 0$, a regular piece which vanishes as $u\rightarrow 0$ and possibly a constant piece. The proposed regularized Yukawa coupling (10) gives the correct result if, after subtracting the entire divergent piece, there is no constant contribution surviving in the limit $u\rightarrow 0$. This is the case for the quintic, where
$\partial_{t}^{3}\mathcal{F}_{0}\big|_{\rm weak}=\sum_{k=0}^{\infty}\frac{a_{(0,k,0)}}{u\,\log^{k+3}u}+{\rm Reg}\,$, (17)
such that the divergent contributions are removed by minimal subtraction. Thus,
$Y_{ttt}^{(0)}=-\lim_{t\rightarrow t_{c}}\left[(\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}-\kappa_{ttt})-{\rm Div}\right]=\kappa_{ttt}\,.$ (18)
where the term $\kappa_{ttt}$ is subtracted explicitly.
3.2 Limit in Kähler moduli space
To see what happens when these divergences are rewritten in terms of $t$, note that the leading singular behavior in (17) reduces to
$\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}(u)=\frac{1}{u\log^{3}u}\left(a_{(0,0,0)}+a_{(0,1,0)}\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$, (19)
with $a_{(0,0,0)}\approx 67.3587$ and $a_{(0,1,0)}\approx 331.301$. In the following, we show that no constant term arises when expressing the expansion (19) in terms of the Kähler modulus $t$. Technically, due to the logarithms this computation is highly involved and we proceed by matching the divergences order by order.
To avoid the complications that arise with complex-valued logarithms, we restrict ourselves to a particular path on the moduli space in which $u$ goes to zero along the positive real axis. In terms of the Kähler modulus this trajectory gets mapped to the positive imaginary axis in the $t$-plane, namely $t=i|t|$. Furthermore, after mapping the branch-cut in the $u$-plane (which we choose to be the standard negative real axis) to the $t$-plane, one can check that the trajectory we have chosen does not cross the branch cut. This is displayed in Figure 1.
Figure 1: Trajectory $u\to 0$ along the real axis and the branch-cut in the $u$-plane, given by the negative real axis, mapped to the $t$-plane using the mirror map (11).
First, we determine the leading order expansion for (19) in terms of $|t|$. For small $u$, the mirror map (11) takes the form
[…]
with the values of $c$ given in Table 1. It follows that
$\log\Delta t=\log u+L+\mathcal{O}\big(\log^{-2}u\big)\,,\qquad\quad L\coloneq\log|\log u|+\log|t_{c}\,c_{(1,0)}|+\frac{c_{(0,1)}}{c_{(1,0)}}\frac{1}{\log u}\,.$ (21)
Even though $\log u<0$ and $c_{(1,0)}<0$, eq.(19) should be a real function so it suffices to take the absolute value inside the logarithms. To match the leading divergence in the limit $u\rightarrow 0$, we evaluate
[…]
which indeed reproduces the leading $1/(u\log^{3}u)$ behavior. However, it also generates two subleading contributions, namely $\log|\log u|/\log u$ and $1/\log u$. The latter is compensated by including an additional term of the form
$\frac{\alpha\ |t_{c}|\,c_{(1,0)}}{\Delta t\log^{3}\Delta t}=\frac{\alpha}{u\log^{3}u}\left(\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$ (23)
on the right hand side. Taking into account eq.(19), by comparison the coefficient $\alpha$ can be determined as
$\alpha=a_{(0,1,0)}+\left(2\log|t_{c}\,c_{(1,0)}|+\frac{c_{(0,1)}}{c_{(1,0)}}\right)a_{(0,0,0)}\,.$ (24)
Finally, we need to cancel the $\log|\log u|$ term. From eq. (21) it follows that this can be achieved from
$\log|\log\Delta t|=\log|\log u|+L\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\,.$ (25)
Indeed, one checks that adding
$2|t_{c}|\,c_{(1,0)}\,a_{(0,0,0)}\frac{\log|\log\Delta t|}{\Delta t\log^{3}\Delta t}=\frac{1}{u\log^{3}u}\left(2a_{(0,0,0)}\frac{\log|\log u|}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$ (26)
exactly cancels the corresponding contribution.
Hence, we have shown that up to order $1/(u\log^{5}u)$, the expansion (19) can be equivalently expressed in terms of $\Delta t$ as
$\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}(\Delta t)=\alpha_{1}\frac{1}{\Delta t\log^{2}\Delta t}+\alpha_{2}\frac{\log|\log\Delta t|}{\Delta t\log^{3}\Delta t}+\alpha_{3}\frac{1}{\Delta t\log^{3}\Delta t}+\mathcal{O}\left(\frac{1}{\Delta t\log^{4}\Delta t}\right)\,$, (27)
with numerical values
$\alpha_{1}=-4.3048\dots\,,\quad\alpha_{2}=-8.6096\dots\,,\quad\alpha_{3}=13.8684\dots\,.$ (28)
It is important to notice that in the limit $u\rightarrow 0$, the terms we have introduced in $\Delta t$ are subleading with respect to the divergences already matched. Specifically, we have the hierarchy of terms
$\frac{1}{\Delta t\log^{2}\Delta t}\gg\frac{\log|\log\Delta t|}{\Delta t\log^{3}\Delta t}\gg\frac{1}{\Delta t\log^{3}\Delta t}\gg\dots\,$, (29)
as $\Delta t$ becomes small. Thus, adding the subleading terms necessary to cancel the $\log|\log u|$ contribution and match the $1/\log u$ piece does not alter the leading divergent behavior that had already been fixed. If one wanted to find the expansion up to $1/\log^{2}u$-order, the same strategy would apply. Higher orders in all the expansions would need to be added, together with the subleading terms $\log|\log\Delta t|/(\Delta t\log^{4}\Delta t)$ and $1/(\Delta t\log^{4}\Delta t)$ which would not affect the previous orders.
In Figure 2 we depict the ratio of the leading divergent terms in (19) and the leading expansion (27) in $\Delta t$ for small values of $u$ (resp. $\Delta t$) to confirm that indeed both expressions are equivalent.
Figure 2: Ratio of eq.(27) and eq.(19) for $u\rightarrow 0$.
Thus, we have shown that indeed no finite constant term is generated in this matching procedure. In other words, performing minimal subtraction on the mirror side does not miss any constant contribution in the original Kähler modulus expansion, leaving us with the correct regularized Yukawa coupling.
We find this quite remarkable, since the structure of the mirror map (11) and prepotential (19) around the conifold point is highly non-trivial. We believe this further supports our proposed regularization procedure.
4 Regularization scheme for the 1-loop prepotential ${\cal F}_{1}$
Another open question is whether this regularization procedure can be extended to higher-genus contributions to the prepotential. As we have seen in eq.(3), similarly to the triple intersection numbers in $\mathcal{F}_{0}$, the genus-one prepotential $\mathcal{F}_{1}$ contains a classical linear term at weak coupling. The question is whether this term can also be obtained via a suitable regularization from the GV Schwinger integral (5).
This integral is divergent, so one has to introduce a regulator, $\epsilon$. Then, for the integral over $s$, we obtain
$\mathcal{F}_{1}=-\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\left(\gamma_{E}+\log(-\epsilon Z_{n})+\mathcal{O}(\epsilon)\right)\,$, (30)
with $\gamma_{E}$ the Euler-Mascheroni constant. There are two different contributions, one coming from $D2$-$D0$-brane bound states and one from pure $D0$-branes. An analogous analysis has been performed in [6] for the non-compact case of the resolved conifold, the only difference being that now one has to perform the sum over the GV invariants so that the two different contributions read
[…]
where $\mu=g_{s}/(e^{\gamma_{E}}\epsilon)$. Note that due to the logarithmic nature of the integral, the regulator appears explicitly when trying to define a regularization procedure for $\mathcal{F}_{1}$. This did not occur for $\mathcal{F}_{0}$, where all the terms involving the regulator $\epsilon$ vanish upon $\zeta$-function regularization. As discussed in [36], operators such as $R^{4}$ in eight-dimensional maximal supergravity or $R^{2}$ in four dimensions with $N=2$ receive logarithmic corrections from massless states. In order to properly account for this divergence, we effectively incorporate a moduli dependent regulator $\epsilon(t)$.
Then, we can get information about the linear term in (3) by taking a derivative with respect to $t$, yielding a function that is analogous to the Yukawa coupling so that we define
$\begin{split}C_{t}&:=\partial_{t}\left(\mathcal{F}_{1}^{D2,D0}+\mathcal{F}_{1}^{D0}\right)\\[3.00003pt] &=-2\pi i\sum_{\beta\neq 0}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\beta\left(\frac{1}{2}+\frac{e^{2\pi i\beta t}}{1-e^{2\pi i\beta t}}\right)-\frac{1}{12}\partial_{t}\log\epsilon(t)\,.\end{split}$ (32)
For the simple choice $\alpha_{0}=1$, $\alpha_{1}=0$, similarly to the case of $Y_{ttt}$, we recover the formal second Chern class of the resolved conifold [33], $c_{2}=-1$. Since the expression already contains the correct world-sheet instanton contributions, Emergence suggests that after regularization, the zero point contribution
$C_{t}^{(0)}:=-\pi i\sum_{\beta\neq 0}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\beta\,\bigg|_{\rm Reg}-\frac{1}{12}\partial_{t}\log\epsilon(t)\,$ (33)
is equal to $(2\pi i)c_{2}/24$. Carrying out the analogous heuristic computation as in the last section motivates the regularization
$C_{t}^{(0)}=-(2\pi i)\lim_{t\to t_{c}}\left[\left(\left.\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}-\frac{c_{2}}{24}\right)-{\rm Div}+\frac{1}{24\pi i}\partial_{t}\log\epsilon(t)\right]\,.$ (34)
Similarly to the case of the genus zero prepotential, this computation is to be done in the strongly coupled M-theory limit, where we need to integrate out the full tower of light states. Of course, the tools to carry this computation are not yet available, but we can use the result from the weakly coupled regime using mirror symmetry. The main difference with respect to $\mathcal{F}_{0}$ is that due to the logarithmic divergence of the initial Schwinger integral, now there is some freedom in choosing the regulator.
4.1 Limit in complex structure moduli space
Let us again focus on the Quintic, where we can rely on the work of [37, 38] to compute $\mathcal{F}_{1}$ [^2] For higher genus computations and related work see e.g. [39, 40, 41]. in terms of the period data of the CY that we already have available. Let $K(t,\bar{t})$ be the Kähler potential for the metric $G_{t\bar{t}}=\partial_{t}\bar{\partial}_{\bar{t}}K(t,\bar{t})$, which can be computed from the symplectic basis of periods in the standard way
$e^{-K}=-i\Pi^{\dagger}\,\Sigma\,\Pi\,,\qquad\qquad\Sigma=\begin{pmatrix}0&0&0&1\\ 0&0&1&0\\ 0&-1&0&0\\ -1&0&0&0\\ \end{pmatrix}\,$. (35)
Then, $\mathcal{F}_{1}$ was determined from a topological string amplitude computation in [37] as
$\mathcal{F}_{1}=\frac{1}{2}\log\left[e^{\left(3+h_{1,1}-\frac{\chi}{12}\right)K}\det(G^{-1})|f|^{2}\right]\,$, (36)
where $f$ is a holomorphic function, the holomorphic anomaly, which should be determined by imposing the appropriate behavior at the singular points of the moduli space. Focusing now on the mirror quintic, the LCS behavior $\mathcal{F}_{1}\rightarrow\frac{50}{24}(t+\bar{t})$, together with imposing regularity at the orbifold point $\psi=0$, fixes $f$ to be
$f=\psi^{\frac{62}{3}}(1-\psi^{-5})^{-\frac{1}{6}}\,.$ (37)
In the topological limit $\bar{t}\rightarrow-i\infty$, the first derivative of the genus one prepotential reads
$\partial_{t}\mathcal{F}_{1}=\frac{\partial\psi}{\partial t}\partial_{\psi}\log\left[e^{\frac{31}{3}K}\det\left(G_{\psi\bar{\psi}}^{-1}\right)^{\frac{1}{2}}\psi^{\frac{62}{6}}(1-\psi^{-5})^{-\frac{1}{12}}\right]\,.$ (38)
This expression holds everywhere in moduli space, which allows us to verify our regularization procedure when approaching the conifold. Working again in the coordinate $u$, for small values we find the expansion
$\left.\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}=\frac{1}{u\log u}\left(\sum_{n,k=0}^{\infty}\frac{(u\log u)^{n}}{\log^{k}u}\tilde{a}_{(n,k)}(u)\right)\,$, (39)
where as usual $\tilde{a}_{(n,k)}(u)$ denotes a polynomial in $u$. For $n=0$, all the constant terms in $\tilde{a}_{(0,k)}(u)$ lead to singular terms which we minimally subtract. Particularly interesting is the term with $n=1$, $k=0$, as it leads to a potential constant contribution, concretely to
$\tilde{a}_{(1,0)}(u)=-0.0219561+\mathcal{O}(u)\,.$ (40)
At first sight, such a constant contribution seems to spoil our proposed regularization. However, just for ${\cal F}_{1}$ there is this extra contribution from the choice of regulator $\epsilon(t)$, which we can choose appropriately in order to cancel the constant term that appears in (34). To make this concrete, in hindsight we choose the regulator to be of the form $\epsilon(t)=\Delta t\,\epsilon$, where $\Delta t$ is defined as in (20) [^3] Note that in the limit $t\to t_{c}$ this regulator is well defined since $\Delta t=|t-t_{c}|>0$.. Then, the regulator dependent term in (34) reads
$\lim_{t\to t_{c}}\frac{1}{24\pi i}\partial_{t}\log\epsilon(t)\approx\lim_{u\to 0}\frac{1}{24\pi i}\frac{du}{dt}\partial_{u}\left[\log u+\log\!|\!\log u|+\log|t_{c}\,c_{(1,0)}|+\frac{c_{(0,1)}}{c_{(1,0)}}\frac{1}{\log u}\right]\,$, (41)
where we have plugged the expansion (21) at leading order. We take the derivative with respect to $u$ and note that inverting the Jacobian in (14) leads to the series expansion
$\frac{du}{dt}=d_{1}(u)+\sum_{n=2}^{\infty}(u\log u)^{n}d_{n}(u)\,$, (42)
where the $d_{n}$ are some infinite series in $u$ with $d_{1}(u)\approx i1.65545u+\mathcal{O}(u^{2})$ being the inverse expansion of (15). Finally, plugging everything in (41) yields
$\lim_{t\to t_{c}}\frac{1}{24\pi i}\partial_{t}\log\epsilon(t)\approx\lim_{u\to 0}\frac{1}{24\pi i}(i1.65545+\mathrm{Reg.})\approx 0.0219561\,.$ (43)
Moreover, it is a straightforward computation to convince oneself that no new divergences are introduced in (34) by choosing this regulator.
In summary, combining the expansion (39) with this regulator in (34) leads to a cancellation of the constant term and therefore yields precisely $c_{2}^{(0)}=(2\pi i)c_{2}/24$ after performing minimal subtraction. As in other cases with such logarithmic divergences, due to inevitable regulator dependence one looses the predictivity of the formalism. Thus, the main observation is that precisely for the case of ${\cal F}_{1}$, where a constant term could emerge, the regulator appears and helps canceling it.
4.2 Limit in Kähler moduli space
Finally, we can do a similar analysis to that of Section 3 and check whether performing the minimal subtraction in terms of $u$ instead of the original Kähler modulus $t$ misses any possible constant contribution. The aim is to find the leading divergence expansion of (39) in terms of $t$. We content ourselves to work up to $1/(u\log^{2}u)$-order, i.e.
$\partial_{t}\mathcal{F}_{1}\bigr|_{\rm weak}(u)=\frac{1}{u\log u}\left(\tilde{a}_{(0,0,0)}+\tilde{a}_{(0,1,0)}\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$, (44)
with $\tilde{a}_{(0,0,0)}=-0.207529$ and $\tilde{a}_{(0,1,0)}=-0.340241$. We proceed in an analogous manner to the case of $\mathcal{F}_{0}$. Recalling the expansion (20), it is clear that
$\frac{|t_{c}|\,c_{(1,0)}\,\tilde{a}_{(0,0,0)}}{\Delta t}=\frac{1}{u\log u}\left(\tilde{a}_{(0,0,0)}-\tilde{a}_{(0,0,0)}\frac{c_{(0,1)}}{c_{(1,0)}}\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$ (45)
matches the leading divergence. In order to match the $1/(u\log^{2}u)$ divergence, we expand $\Delta t\log\Delta t$ and find
$\frac{|t_{c}|c_{(1,0)}\,\beta}{\Delta t\log\Delta t}=\frac{\beta}{u\log u}\left(\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,.$ (46)
By setting
$\beta=\tilde{a}_{(0,1,0)}+\tilde{a}_{(0,0,0)}\frac{c_{(0,1)}}{c_{(1,0)}}\,$, (47)
the expansion (44) can equivalently be expressed in terms of $t$ as
$\partial_{t}\mathcal{F}_{1}(\Delta t)=\alpha_{1}\frac{1}{\Delta t}+\alpha_{2}\frac{1}{\Delta t\log\Delta t}+\mathcal{O}\left(\frac{1}{\Delta t\log^{2}\Delta t}\right)$ (48)
with $\alpha_{1}=|t_{c}|c_{(1,0)}\,\tilde{a}_{(0,0,0)}\approx 0.0132$ and $\alpha_{2}=|t_{c}|c_{1,0}\,\beta\approx-0.0132$. Note that this is in accordance with the expected behavior of $\mathcal{F}_{1}$ around the conifold, namely
$2\pi i\,\mathcal{F}_{1}\sim-\frac{1}{12}\log(t-t_{c})\,$, (49)
as e.g. found in [42].
In Figure 3 we show the ratio of both expansions to confirm that around $u\to 0$ both expressions are equivalent.
Figure 3: Ratio of eq.(48) and eq.(44) for $u\rightarrow 0$.
Again, no finite constant appears when inverting the mirror map or, differently said, performing the minimal subtraction on the mirror side, which is arguably easier, does not miss any possible constant contribution in the original Kähler expansion. Once more, we remark that the fact that the minimal subtraction does not depend on whether we consider the Kähler or the mirror complex structure moduli space provides non-trivial support for our proposed regularization of the sum over the GV invariants.
5 Conclusion
Motivated by questions raised during talks given on the original paper [32], in this article we have extended the emergence computation to the linear term in the 1-loop prepotential ${\cal F}_{1}$. In contrast to ${\cal F}_{0}$, here the result is less predictive as the regulator plays an important role. However, it is still not in contradiction with the proposed regularization, as precisely for the case where a potential constant could appear, the appearing regulator can cancel it. We have shown this in detail for the Quintic threefold and expect this computation to be analogous for the other $14$ CY threefolds with $h^{1,1}=1$. In addition, for both ${\cal F}_{0}$ and ${\cal F}_{1}$, we have shown that performing the limit of the regularization procedure in the more tractable complex structure moduli space of the mirror CY ultimately yields the same result as taking the limit in the physical Kähler moduli space. This a postereori confirms the method used in [32].
It would also be interesting to gain a better understanding of the regularization procedure for CY manifolds with multiple Kähler moduli. The methods to deal with such cases were introduced in [32] and two CYs with $h_{11}=2$ were discussed in detail, with the final results being encouraging but not yet fully conclusive. That’s because new issues, like the choice of the codimension two singular point in the Kähler moduli space and the correct path towards, which remain to be explored in more detail. We hope to return to these in the future.
Nevertheless, we think it is fair to say that by now we have promising evidence for the validity of the M-theoretic Emergence Proposal, at least for $1/2$-BPS saturated amplitudes. For the general non-BPS amplitudes it is expected that one really needs a fully fledged QG theory of M-theory. From this broader perspective, the emerging nature of $1/2$-BPS amplitudes is just a reflection of the emerging nature of interactions in quantum M-theory. The expectation is that, analogous to the BFSS matrix theory, an ordinary commuting space-time emerges as a particular solution to the equations of motion of such a theory. Interactions, including gravity, then arise from quantum effects, potentially through the evaluation of appropriate Schwinger integrals over the spectrum of quantum M-theory.
Acknowledgments.
We are grateful to Thorsten Schimannek and Niccolò Cribiori for discussions on the emergence of the one-loop prepotential. We also thank Antonia Paraskevopoulou for helpful comments on the draft. R.B. acknowledges the hospitality of the Corfu Summer Institute 2025, where part of the research presented here has been initiated.

### 31 Mar 2026 v2

@@
 After reviewing the current status of this proposal, we focus on four-dimensional compactifications with $N=2$ supersymmetry, where kinetic terms are encoded in topological string amplitudes, such as the prepotential $\mathcal{F}_{0}$.
-Evidence for the emergence of such terms was provided recently, where in particular it was shown that the classical cubic term in $\mathcal{F}_{0}$ can be obtained by integrating out the light towers of states in the M-theory limit, using a novel regularization of the infinite sum over Gopakumar/Vafa invariants.
+Evidence for the emergence of such terms was provided recently, where in particular it was shown that the classical cubic term in ${\cal F}_{0}$ can be obtained by integrating out the light towers of states in the M-theory limit, using a novel regularization of the infinite sum over Gopakumar/Vafa invariants.
 We address two issues that were left open.
@@
 Similarly to the triple intersection number in $\mathcal{F}_{0}$, the genus-one prepotential $\mathcal{F}_{1}$ contains a classical term at weak coupling, which for a one-parameter CY threefold reads
-$\mathcal{F}_{1}(t)={2\pi i}\left(\frac{c_{2}}{24}t+\frac{1}{2\pi i}\sum_{\beta\in H_{2}(X,\mathbb{Z})}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right){\rm Li}_{1}(e^{2\pi i\beta\cdot t})\right)\,$, (3)
+$\mathcal{F}_{1}(t)={2\pi i}\ \left(\frac{c_{2}}{24}t-\frac{1}{2\pi i}\sum_{\beta\in H_{2}(X,\mathbb{Z})}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right){\rm Li}_{1}(e^{2\pi i\beta\cdot t})\right)\,$, (3)
 where $c_{2}$ is the second Chern class of the manifold and $\alpha_{1}^{\beta}$ are the genus one GV invariants [30, 31].
@@
 where $Z_{\beta,n}=\frac{2\pi i}{g_{s}}(\beta\cdot t-n)$ is the central charge of the supersymmetry algebra. Similarly, for $\mathcal{F}_{1}$ Gopakumar/Vafa obtained
-$\mathcal{F}_{1}(t)=\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\int_{0}^{\infty}\frac{ds}{s}e^{sZ_{\beta,n}}\,.$ (5)
+$\mathcal{F}_{1}(t)=-\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\int_{0}^{\infty}\frac{ds}{s}e^{sZ_{\beta,n}}\,.$ (5)
 We recall that, after employing a minimal subtraction of the UV divergence at $t\to 0$ and subsequently applying $\zeta$-function regularization, the complete $\mathcal{F}_{0}$ and $\mathcal{F}_{1}$ for the (non-compact) resolved conifold have been reproduced in [6].
@@
 Implementing the zeta function regularization $\sum_{k=1}^{\infty}1=\zeta(0)=-1/2$ one writes
-$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{n}n^{3}=-\sum_{n,k=1}^{\infty}\alpha_{n}n^{3}=-\lim_{\tau\to 0}\sum_{n,k=1}^{\infty}\alpha_{n}n^{3}e^{-2\pi nk\tau}=-\lim_{\tau\to 0}\sum_{n=1}^{\infty}\alpha_{n}n^{3}\frac{q^{n}}{1-q^{n}}\end{split}$ (7)
+$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{0}^{n}n^{3}=-\sum_{n,k=1}^{\infty}\alpha_{0}^{n}n^{3}=-\lim_{\tau\to 0}\sum_{n,k=1}^{\infty}\alpha_{0}^{n}n^{3}e^{-2\pi nk\tau}=-\lim_{\tau\to 0}\sum_{n=1}^{\infty}\alpha_{0}^{n}n^{3}\frac{q^{n}}{1-q^{n}}\end{split}$ (7)
 with $q=\exp(-2\pi\tau)$.
@@
 Hence, setting for ease of notation $(2\pi i)^{3}/g_{s}^{2}=1$, we can express the last term in (7) as
-$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{n}n^{3}=-\lim_{\tau\to 0}\sum_{n}^{\infty}\alpha_{n}n^{3}\frac{q^{n}}{1-q^{n}}=-\lim_{t\to 0}\left[(\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}-\kappa_{ttt})\right]\end{split}$ (8)
+$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{0}^{n}n^{3}=-\lim_{\tau\to 0}\sum_{n}^{\infty}\alpha_{0}^{n}n^{3}\frac{q^{n}}{1-q^{n}}=-\lim_{t\to 0}\left[(\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}-\kappa_{ttt})\right]\end{split}$ (8)
 where we have identified $t=i\tau$ and $\mathcal{F}_{0}|_{\rm weak}$ denotes the prepotential that one determines by determining periods in the complex structure moduli and applying the mirror map.
@@
 Thus, adding the subleading terms necessary to cancel the $\log|\log u|$ contribution and match the $1/\log u$ piece does not alter the leading divergent behavior that had already been fixed.
-If one wanted to find the expansion up to $1/\log^{2}u$-order, the same strategy would apply.
+If one wanted to find the expansion of (19) up to order $1/(u\log^{5}u)$, the same strategy would apply.
 Higher orders in all the expansions would need to be added, together with the subleading terms $\log|\log\Delta t|/(\Delta t\log^{4}\Delta t)$ and $1/(\Delta t\log^{4}\Delta t)$ which would not affect the previous orders.
@@
 This integral is divergent, so one has to introduce a regulator, $\epsilon$. Then, for the integral over $s$, we obtain
-$\mathcal{F}_{1}=-\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\left(\gamma_{E}+\log(-\epsilon Z_{n})+\mathcal{O}(\epsilon)\right)\,$, (30)
+$\mathcal{F}_{1}=\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\left(\gamma_{E}+\log(-\epsilon Z_{\beta,n})+\mathcal{O}(\epsilon)\right)\,$, (30)
 with $\gamma_{E}$ the Euler-Mascheroni constant.
@@
 An analogous analysis has been performed in [6] for the non-compact case of the resolved conifold, the only difference being that now one has to perform the sum over the GV invariants so that the two different contributions read
-[…]
+$\begin{split}\mathcal{F}_{1}^{D0,D2}&=\sum_{\beta\neq 0}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\left(-\pi i\,\beta t+\log\big(1-e^{2\pi i\beta t}\big)\right)\,,\\[2.84544pt] \mathcal{F}_{1}^{D0}&=\frac{1}{12}\log\mu\,.\end{split}$ (31)
 where $\mu=g_{s}/(e^{\gamma_{E}}\epsilon)$.
@@
 is equal to $(2\pi i)c_{2}/24$. Carrying out the analogous heuristic computation as in the last section motivates the regularization
-$C_{t}^{(0)}=-(2\pi i)\lim_{t\to t_{c}}\left[\left(\left.\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}-\frac{c_{2}}{24}\right)-{\rm Div}+\frac{1}{24\pi i}\partial_{t}\log\epsilon(t)\right]\,.$ (34)
+$C_{t}^{(0)}=-(2\pi i)\lim_{t\to t_{c}}\left[\left(\left.\frac{1}{2\pi i}\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}-\frac{c_{2}}{24}\right)-{\rm Div}+\frac{1}{24\pi i}\partial_{t}\log\epsilon(t)\right]\,.$ (34)
 Similarly to the case of the genus zero prepotential, this computation is to be done in the strongly coupled M-theory limit, where we need to integrate out the full tower of light states.
@@
 Then, $\mathcal{F}_{1}$ was determined from a topological string amplitude computation in [37] as
-$\mathcal{F}_{1}=\frac{1}{2}\log\left[e^{\left(3+h_{1,1}-\frac{\chi}{12}\right)K}\det(G^{-1})|f|^{2}\right]\,$, (36)
+$\mathcal{F}_{1}=\pi i\log\left[e^{\left(3+h_{1,1}-\frac{\chi}{12}\right)K}\det(G^{-1})|f|^{2}\right]\,$, (36)
 where $f$ is a holomorphic function, the holomorphic anomaly, which should be determined by imposing the appropriate behavior at the singular points of the moduli space.
-Focusing now on the mirror quintic, the LCS behavior $\mathcal{F}_{1}\rightarrow\frac{50}{24}(t+\bar{t})$, together with imposing regularity at the orbifold point $\psi=0$, fixes $f$ to be
+Focusing now on the mirror quintic, the LCS behavior $\mathcal{F}_{1}\rightarrow 2\pi i\ \frac{50}{24}(t+\bar{t})$, together with imposing regularity at the orbifold point $\psi=0$, fixes $f$ to be
 $f=\psi^{\frac{62}{3}}(1-\psi^{-5})^{-\frac{1}{6}}\,.$ (37)
 In the topological limit $\bar{t}\rightarrow-i\infty$, the first derivative of the genus one prepotential reads
-$\partial_{t}\mathcal{F}_{1}=\frac{\partial\psi}{\partial t}\partial_{\psi}\log\left[e^{\frac{31}{3}K}\det\left(G_{\psi\bar{\psi}}^{-1}\right)^{\frac{1}{2}}\psi^{\frac{62}{6}}(1-\psi^{-5})^{-\frac{1}{12}}\right]\,.$ (38)
+$\partial_{t}\mathcal{F}_{1}=2\pi i\ \frac{\partial\psi}{\partial t}\partial_{\psi}\log\left[e^{\frac{31}{3}K}\det\left(G_{\psi\bar{\psi}}^{-1}\right)^{\frac{1}{2}}\psi^{\frac{62}{6}}(1-\psi^{-5})^{-\frac{1}{12}}\right]\,.$ (38)
 This expression holds everywhere in moduli space, which allows us to verify our regularization procedure when approaching the conifold.
 Working again in the coordinate $u$, for small values we find the expansion
-$\left.\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}=\frac{1}{u\log u}\left(\sum_{n,k=0}^{\infty}\frac{(u\log u)^{n}}{\log^{k}u}\tilde{a}_{(n,k)}(u)\right)\,$, (39)
+$\left.\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}=\frac{2\pi i}{u\log u}\left(\sum_{n,k=0}^{\infty}\frac{(u\log u)^{n}}{\log^{k}u}\tilde{a}_{(n,k)}(u)\right)\,$, (39)
 where as usual $\tilde{a}_{(n,k)}(u)$ denotes a polynomial in $u$.
@@
 We content ourselves to work up to $1/(u\log^{2}u)$-order, i.e.
-$\partial_{t}\mathcal{F}_{1}\bigr|_{\rm weak}(u)=\frac{1}{u\log u}\left(\tilde{a}_{(0,0,0)}+\tilde{a}_{(0,1,0)}\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$, (44)
+$\partial_{t}\mathcal{F}_{1}\bigr|_{\rm weak}(u)=\frac{2\pi i}{u\log u}\left(\tilde{a}_{(0,0,0)}+\tilde{a}_{(0,1,0)}\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$, (44)
 with $\tilde{a}_{(0,0,0)}=-0.207529$ and $\tilde{a}_{(0,1,0)}=-0.340241$. We proceed in an analogous manner to the case of $\mathcal{F}_{0}$. Recalling the expansion (20), it is clear that
@@
 the expansion (44) can equivalently be expressed in terms of $t$ as
-$\partial_{t}\mathcal{F}_{1}(\Delta t)=\alpha_{1}\frac{1}{\Delta t}+\alpha_{2}\frac{1}{\Delta t\log\Delta t}+\mathcal{O}\left(\frac{1}{\Delta t\log^{2}\Delta t}\right)$ (48)
+$\partial_{t}\mathcal{F}_{1}(\Delta t)=\alpha_{1}\frac{2\pi i}{\Delta t}+\alpha_{2}\frac{2\pi i}{\Delta t\log\Delta t}+\mathcal{O}\left(\frac{1}{\Delta t\log^{2}\Delta t}\right)$ (48)
 with $\alpha_{1}=|t_{c}|c_{(1,0)}\,\tilde{a}_{(0,0,0)}\approx 0.0132$ and $\alpha_{2}=|t_{c}|c_{1,0}\,\beta\approx-0.0132$.
 Note that this is in accordance with the expected behavior of $\mathcal{F}_{1}$ around the conifold, namely
-$2\pi i\,\mathcal{F}_{1}\sim-\frac{1}{12}\log(t-t_{c})\,$, (49)
+$\mathcal{F}_{1}\sim-\frac{1}{12}\log(t-t_{c})\,$, (49)
 as e.g. found in [42].